In Newton’s ring experiment the diameter of 16





th bright ring changes from 1.8 cm to 1.5cm when a liquid is





introduced the plate and the lens. Find the refractive index of the liquid and the diameter of the 5





th dark ring

1
Expert's answer
2021-11-09T16:36:20-0500

"d_{b16} =1.8cm"

"d'_{b16} =1.5cm"

"r_{bk}=\sqrt{(k-\frac{1}{2})\frac{\lambda R}{n}}"

"\frac{r_{b16}}{r'_{b16}}=\frac{d_{b16}}{d'_{b16}}=\frac{1.8}{1.5}= 1.2"

"r_{b16}=\sqrt{(16-\frac{1}{2})\frac{\lambda R}{n_1}}"

"r'_{b16}=\sqrt{(16-\frac{1}{2})\frac{\lambda R}{n_2}}\ (1)"

"\frac{r_{b16}}{r'_{b16}}=\sqrt{\frac{n_2}{n_1}}"

"n_1\approx 1 \text{ suppose the expiration is in the air}"

"n_2 =1.2^2=1.44"

"r_{dk}=\sqrt{k\frac{\lambda R}{n}}"

"\text{From the formula (1)}"

"\sqrt{\frac{\lambda R}{n_2}}=r'_{b16}*\sqrt{\frac{1}{15.5}}\text{ then}"

"r'_{d5}=\sqrt{5}*\frac{d'_{b16}}{2}*\sqrt{\frac{1}{15.5}}=0.75*\sqrt{\frac{5}{15.5}}=0.43cm"


"\text{Answer:}"

"1.44 -\text{refractive index of a liquid}"

"0.43cm - \text{radius 5 of the dark ring when immersed in liquid}"